# OLMES: A Standard for Language Model Evaluations

### v1

@@ 3.2 Which few-shot examples to use?
Popularized by Brown et al., (2020), it is customary to provide examples of the task to the model through few-shot examples, as this is an effective and universal way to convey a task to an LLM. For example, the MMLU task (Hendrycks et al.,, 2021) originally came with a fixed 5-shot prompt which is generally used in evaluation (Beeching et al.,, 2023; Gemma Team et al.,, 2024; Jiang et al.,, 2023; Touvron et al., 2023b, ; AI@Meta,, 2024) resulting in a more reproducible results than many other tasks. [^4] Sometimes sampled examples are used also for MMLU (MosaicML,, 2024). Even for MMLU, noticeable discrepancies have been found, due to other differences in prompt formatting (Mai and Liang,, 2024)). For other tasks, both the number of shots and the way in which they are sampled have varied in different evaluation setups. For example, to evaluate on HellaSwag, Beeching et al., (2023) sampled 10-shot whereas HELM Liang et al., (2023) uses 0-shot; within Beeching et al., (2023), a range of 25-shot, 10-shot, 5-shot was sampled for ARC-Challenge, HellaSwag and WinoGrande respectively.
OLMES standardizes a manually curated 5-shots prompt for each task (from its training set), ensuring that the examples are of good quality and cover the label space in a balanced way (e.g., avoiding 4 A’s and 1 B among the 5 answers). [^5] More details on curating the examples can be found in Appendix F. Restricting to 5 in-context examples helps limit computational overhead, similar to HELM (Liang et al.,, 2023). Analysis suggests that going beyond 5 shots generally does not provide meaningful differences in scores (Brown et al.,, 2020; Barton,, 2024). These manually curated shots for each task can be downloaded from https://allenai.org/data/olmes.
@@ 3.3 How to normalize LLM probabilities for CF?
When using the cloze/completion formulation (CF) for multiple-choice questions, the LLM returns $P(a_{i}|q)$ to denote the probability for an answer choice $a_{i}$ given a question prompt $q$. Ranking solely based on the probability may heavily favor shorter answers with fewer tokens. To work around this issue, different normalization methods have been used in the literature, which we categorize below:
@@
  • token: $\ln(P(a_{i}|q))/\operatorname{num\_tokens}(a_{i})$, which normalizes the log-probability by the number of tokens in the answer (Brown et al.,, 2020)
@@
  • character: $\ln(P(a_{i}|q))/\operatorname{num\_characters}(a_{i})$, which normalizes the log-probability by the number of characters in the answer, used by Llama models (Touvron et al., 2023a, ) and Eleuther AI LM Harness (Gao et al.,, 2023; Biderman et al.,, 2024).
@@
  • pmi: $\ln(P(a_{i}|q)/P(a_{i}|u))$ where $u=$"Answer:" is the unconditional prompt, which normalizes by dividing by LLM probability of the same answer string without the presence of the question (like "Answer: <answer_string>"). This can be considered a form of pointwise-mutual-information (PMI) and was explored further in other works (Holtzman et al.,, 2021).
@@
Efforts like Liang et al., (2023); Gao et al., (2023); Biderman et al., (2024) compare and support comparisons of different normalization approaches, leaving it an open question as to how to make a decision. See Appendix C.2 for further discussions around different normalizations.
@@
To choose a normalization scheme in OLMES, we evaluate the models on each dataset using CF, comparing the effect of the 4 normalization techniques. Table 3 shows for each task, how often each normalization is empirically the best across the 15 models. Detailed scores per model are in Appendix C.2.
OLMES specifies the “pmi” normalization for ARC-Challenge, CommonsenseQA, and OpenbookQA. The answer choices in these datasets tend to contain unexpected words or phrases that are less likely for models to generate (e.g., “Whirlpool bath” compared to “Bathtub”). The pmi normalization adjusts for this by taking into account the a priori likelihood of the answers. This is consistent with other findings (Holtzman et al.,, 2021) and some existing evaluation practices, e.g., Brown et al., (2020) selectively uses this normalization for ARC and OpenbookQA, and Touvron et al., 2023a ; Touvron et al., 2023b for OpenbookQA. Computing the extra unconditional likelihood incurs some computation overhead, thus OLMES avoids this normalization for other datasets where there is no strong empirical or theoretical reason to choose this approach.
OLMES specifies the “character” normalization for ARC-Easy, HellaSwag, PIQA, Social IQa and MMLU. Based on our experiments, it is empirically the normalization technique that gives the best scores [^6] Tie for PIQA, and second-best for MMLU. for these datasets, and less computationally expensive than the “pmi” normalization. It also has the advantage (unlike the “token” normalization) of already being implemented (as acc_norm) in the Eleuther LM Evaluation Harness, where it is generally available for multiple-choice tasks (Gao et al.,, 2023; Gao,, 2021). It is also used in the Hugging Face Open LLM Leaderboard (Beeching et al.,, 2023) for ARC-Challenge and HellaSwag, in Touvron et al., 2023a ; Touvron et al., 2023b ’s evaluations as the default, (with select datasets as exceptions), as well as reported in various works like Biderman et al., (2023); Almazrouei et al., (2023).
In general, we observe little difference between the OLMES recommendation and the empirically best (“oracle”) normalization for each task and model, see “diff best” column in Table 3 (Table 8 in Appendix C.2 has more details).
3.4 What task formulation to use, MCF or CF?
As LLMs have gotten stronger, the MCQA task formats have gradually changed from CF to MCF. For instance, ARC-Challenge is often evaluated using the CF approach (Touvron et al., 2023a, ; Touvron et al., 2023b, ; Almazrouei et al.,, 2023; Beeching et al.,, 2023), but has switched to MCF for stronger models like OpenAI, (2024); AI@Meta, (2024), appearing with identical names like “25-shot ARC-Challenge”. As an example, AI@Meta, (2024) reports a 25-shot MCF ARC-Challenge score for the Llama-3 8B model of 78.6% vs 59.2% for the 25-shot CF on the Hugging Face Open LLM Leaderboard. As performance on a multiple-choice task gets closer to 100%, the CF approach lags behind and give significantly less signal about a model’s actual performance. On the other hand, MMLU is almost exclusively evaluated using the MCF approach, which often results in near-random performance for weaker models (Beeching et al.,, 2023).
In OLMES, we argue that the CF formulation provides a useful evaluation of task knowledge for models that have not yet acquired the skill of answering multiple-choice questions using MCF. On the other hand, MCF is a more realistic formulation for models that can “understand” this format, yielding higher and more representative scores (Robinson et al.,, 2023; OpenAI,, 2024). See Appendix C.1 for further discussion.
Figure 1: Performance on MMLU validation set during the training of OLMo-1.7-7B model. During early training, there is good signal from CF while MCF is random. Around 400B tokens, the model starts gaining the ability on the MCF format, becoming a stronger signal than CF.
Table 3: Summary of CF normalization comparisons. “win percentage” shows how often each normalization was best across the 15 models. “diff best” shows that there is in general minimal difference between the OLMES normalization and the empirically optimal normalization for each task (difference out of 100%).
@@ 3.4 What task formulation to use, MCF or CF?
task none char tok pmi best OLMES
@@
We can see an explicit example of a model acquiring the “understanding” of MCF during training in Figure 1, showing the OLMo-1.7-7B model (Groeneveld et al.,, 2024; AI2 blog,, 2024) evaluated on the MMLU validation set in both CF and MCF variations. The plot suggests that model starts learning the MCF task format after about 400 billion training tokens, so in early training CF provides a better signal, while MCF is significantly better in late training.
To further study this phenomenon, we evaluate the CF and MCF versions on each task and model. [^7] More detailed numbers can be found in Tables 5 and 6 in the Appendix B Figure 2 shows for each task, the CF and MCF performance for the 15 models ordered along the x-axis by overall performance on all tasks. For instance, on ARC-Challenge, we see a clear distinction where the weakest 8 models have near-random performance on the MCF version of the task, yet above random when using CF which offers a better signal to the relative strength of models. For the stronger models, the MCF version clearly outscores the CF version, and is a much better representation of task performance compared to the flatter trends using CF (for Llama3-70B the MCF score is 93.7% (6.3% error) while the CF score is just 69.0% (31% error), a nearly 5x difference in error rate!).
@@ 4 OLMES: Summary and results
  • Use test set when available, otherwise validation. Sample 1000 instances if more than 1500
@@
  • Follow recommendations for all other evaluation details (in Appendix A)
Table 4 reports scores for different models on popular benchmark tasks, following our evaluation standard, OLMES.
@@
OLMo-1.7-7B $66.9^{\dagger}$ $83.6^{\dagger}$ 85.9 $85.8^{\dagger}$ 80.1 $54.4^{\dagger}$ $68.6^{\dagger}$ 80.3 $76.1^{\dagger}$ 73.6 75.5
@@ 5 Related work
With model releases, performance on popular datasets is used to gauge the progress achieved e.g., OpenAI, (2024)t showing superhuman performance on benchmarks like MMLU. Such evaluation also guides the community efforts towards understanding and sharing findings on what it takes to build a strong model (Touvron et al., 2023a ; Touvron et al., 2023b ; Biderman et al., (2023); Almazrouei et al., (2023); MosaicML, (2023); Jiang et al., (2023); Gemma Team et al., (2024); Groeneveld et al., (2024) inter alia).
However, given a model and a dataset, even for the frequently used datasets, there are varied practices in how accuracy on them is measured. Various work has shown that model evaluations are vulnerable to differences such as option position changes in multiple-choice questions (Zheng et al.,, 2024; Li et al.,, 2024), choice symbols, re-ordering of answer options, changing number of answer options (Wang et al.,, 2024), and task formulation (Alzahrani et al.,, 2024; Robinson et al.,, 2023; Khatun and Brown,, 2024; Wiegreffe et al.,, 2023). Even minor formatting changes can cause large, generally arbitrary, score variations (Sclar et al.,, 2023). The Holistic Evaluation of Language Models (HELM) benchmark (Liang et al.,, 2023), the Hugging Face Open LLM Leaderboard (Beeching et al.,, 2023), Mosaic Eval Gauntlet (Barton,, 2024) and Eleuther AI LM Harness (Gao et al.,, 2023; Biderman et al.,, 2024) present efforts toward greater transparency and reproducibility of LM evaluations. Liang et al., (2023); Gao et al., (2023); Biderman et al., (2024) describe and provide support for variability factors, presenting them as open choices to researchers and users. Even when specifications are made, the rationale is not always documented and thus not followed by others in subsequent work (see Table 1).
@@ 6 Discussion
By identifying and reviewing common evaluation practices in the community, and performing experiments to investigate open questions, we present OLMES – an open, documented, reproducible, and practical evaluation standard.
@@
Future work and limitations. Our efforts for future work include adding more tasks to OLMES, covering tasks beyond MCQA such as generative tasks and chain-of-thought prompting. This encompasses standardizing how answers are extracted for evaluation, and for chat models how to split the prompt into messages. We also encourage the community to contribute to OLMES, extending the principles of OLMES to new tasks.
@@
We hope OLMES will be a meaningful step towards standardizing evaluations, and incorporated into evaluation code bases for broad usage. This will facilitate robust and simplified comparisons of model performances, both for researchers during model training and development and for developers choosing models to build upon.
Acknowledgments and Disclosure of Funding

### v2

@@ 3.2 Which few-shot examples to use?
Popularized by Brown et al. (2020), it is customary to provide examples of the task to the model through few-shot examples, as this is an effective and universal way to convey a task to an LLM. For example, the MMLU task (Hendrycks et al., 2021) originally came with a fixed 5-shot prompt which is generally used in evaluation (Beeching et al., 2023; Gemma Team et al., 2024; Jiang et al., 2023; Touvron et al., 2023b; AI@Meta, 2024) resulting in more reproducible results than many other tasks. [^4] Sometimes sampled examples are used also for MMLU (MosaicML, 2024). Even for MMLU, noticeable discrepancies have been found, due to other differences in prompt formatting (Mai and Liang, 2024). For other tasks, both the number of shots and the way in which they are sampled have varied in different evaluation setups. For example, to evaluate on HellaSwag, Beeching et al. (2023) sampled 10-shot whereas HELM (Liang et al., 2023) uses 0-shot; within Beeching et al. (2023), a range of 25-shot, 10-shot, 5-shot was sampled for ARC-Challenge, HellaSwag and WinoGrande respectively.
OLMES standardizes a manually curated 5-shots prompt for each task (from its training set), ensuring that the examples are of good quality and cover the label space in a balanced way (e.g., avoiding 4 A’s and 1 B among the 5 answers). [^5] More details on curating the examples can be found in Appendix G. Restricting to 5 in-context examples helps limit computational overhead, similar to HELM (Liang et al., 2023). Analysis suggests that going beyond 5 shots generally does not provide meaningful differences in scores (Brown et al., 2020; Barton, 2024). The manually curated shots for each task can be downloaded from https://github.com/allenai/olmes.
@@ 3.3 How to normalize LLM probabilities for CF?
When using the completion/cloze formulation (CF) for multiple-choice questions, the LLM returns $P(a_{i}|q)$, the probability for an answer choice $a_{i}$ given a question prompt $q$. Ranking solely based on the probability may heavily favor shorter answers with fewer tokens. To work around this issue, different normalization methods have been used in the literature, which we categorize below:
@@
  • token: $\ln(P(a_{i}|q))/\operatorname{num\_tokens}(a_{i})$, which normalizes the log-probability by the number of tokens in the answer (Brown et al., 2020).
@@
  • character: $\ln(P(a_{i}|q))/\operatorname{num\_characters}(a_{i})$, which normalizes the log-probability by the number of characters in the answer, used by Llama models (Touvron et al., 2023a) and Eleuther AI LM Harness (Gao et al., 2023; Biderman et al., 2024).
@@
  • pmi: $\ln(P(a_{i}|q)/P(a_{i}|u))$ where $u=$"Answer:" is an unconditional prompt, which normalizes by dividing by the LLM probability of the same answer string without the presence of the question. This can be considered a form of pointwise-mutual-information (PMI) and was explored further in other works (Holtzman et al., 2021).
@@
Efforts like Liang et al. (2023); Gao et al. (2023); Biderman et al. (2024) compare and support comparisons of different normalization approaches, leaving it an open question as to how to make a decision. See Appendix C.2 for further discussions around different normalizations.
@@
To choose a normalization scheme in OLMES, we evaluate the models on each dataset, comparing the effect of the 4 normalization techniques. Table 3 shows for each task, how often each normalization is empirically the best across the 15 models. Detailed scores per model are in Appendix C.2.
@@
task none char tok pmi oracle OLMES
@@
Table 3: Summary of CF normalization comparisons. “win percentage” shows how often each normalization was best across the 15 models. “diff oracle” (difference between the OLMES recommendation and the empirically best normalization for each task and model) shows that there is in general minimal difference between the OLMES normalization and the oracle optimal normalization for each task (difference out of 100%).
OLMES specifies the “pmi” normalization for ARC-Challenge, CommonsenseQA, and OpenbookQA. The answer choices in these datasets tend to contain unexpected words or phrases that are less likely for models to generate (e.g., “Whirlpool bath” compared to “Bathtub”). The pmi normalization adjusts for this by taking into account the a priori likelihood of the answers. This is consistent with other findings (Holtzman et al., 2021) and some existing evaluation practices, e.g., Brown et al. (2020) selectively uses this normalization for ARC and OpenbookQA, and Touvron et al. (2023a, b) for OpenbookQA. Computing the extra unconditional likelihood incurs some computation overhead, thus OLMES avoids this normalization for other datasets where there is no strong empirical or theoretical reason to choose this approach.
OLMES specifies the “character” normalization for ARC-Easy, HellaSwag, PIQA, Social IQa and MMLU. Based on our experiments, it is empirically the normalization technique that gives the best scores [^6] Tie for PIQA, and second-best for MMLU. for these datasets, and less computationally expensive than the “pmi” normalization. It also has the advantage (unlike the “token” normalization) of already being implemented (as acc_norm) in the Eleuther LM Evaluation Harness, where it is generally available for multiple-choice tasks (Gao et al., 2023; Gao, 2021). It is also used in the Hugging Face Open LLM Leaderboard (Beeching et al., 2023) for ARC-Challenge and HellaSwag, in Touvron et al. (2023a, b)’s evaluations as the default, (with select datasets as exceptions), as well as reported in various works like Biderman et al. (2023); Almazrouei et al. (2023).
In general, we observe little difference between the OLMES recommendation and the empirically best (“oracle”) normalization for each task and model, see “diff oracle” column in Table 3 (Table 9 in Appendix C.2 has more details).
3.4 What task formulation to use, MCF or CF?
As LLMs have gotten stronger, the MCQA task formats have gradually changed from CF to MCF. For instance, ARC-Challenge was often evaluated using the CF approach (Touvron et al., 2023a, b; Almazrouei et al., 2023; Beeching et al., 2023), but has switched to MCF for stronger models like OpenAI (2024); AI@Meta (2024), appearing with identical names like “25-shot ARC-Challenge”. As an example, AI@Meta (2024) reports a 25-shot MCF ARC-Challenge score for the Llama-3 8B model of 78.6% vs 60.2% for the 25-shot CF on the Hugging Face Open LLM Leaderboard. As performance on a multiple-choice task gets closer to 100%, the CF approach lags behind due to its inherent limitations, giving significantly less signal about a model’s actual performance. On the other hand, MMLU is almost exclusively evaluated using the MCF approach, which often results in near-random performance for weaker models (Beeching et al., 2023).
In OLMES, we argue that the CF formulation provides a useful evaluation of task knowledge for models that have not yet acquired the skill of answering multiple-choice questions using MCF. On the other hand, MCF is a more realistic formulation for models that can “understand” this format, yielding higher and more representative scores (Robinson et al., 2023; OpenAI, 2024). See Appendix C.1 for further discussion.
Figure 1: Performance on MMLU validation set during the training of OLMo-7B-0424 model. During early training, there is good signal from CF while MCF is random. Around 400B tokens, the model starts gaining the ability on the MCF format, becoming a stronger signal than CF.
We can see an explicit example of a model acquiring the “understanding” of MCF during training in Figure 1, showing the OLMo-7B-0424 model (Groeneveld et al., 2024; AI2 blog, 2024) evaluated on the MMLU validation set in both CF and MCF variations. The plot suggests that model starts learning the MCF task format after about 400 billion training tokens, so in early training CF provides a better signal, while MCF is significantly better in late training where CF levels off.
@@ 3.4 What task formulation to use, MCF or CF?
To further study this phenomenon, we evaluate the CF and MCF versions for each task and model. [^7] More detailed numbers can be found in Tables 6 and 7 in the Appendix B. Figure 2 shows for each task, the MCF and CF performances for the 15 models ordered along the x-axis by overall performance on all tasks. For instance, on ARC-Challenge, we see a clear distinction where the weakest 8 models have near-random performance on the MCF version of the task, yet above random when using CF which offers a better signal to the relative strength of models. For the stronger models, the MCF version clearly outscores the CF version, and is a much better representation of task performance compared to the flatter trends using CF (for Llama3-70B the MCF score is 93.7% (6.3% error) while the CF score is just 69.0% (31% error), a nearly 5x difference in error rate!).
@@
3.5 Other implementation details
  • For MMLU: use macro average (over 57 tasks) rather than micro average (over 14042 instances), following AI@Meta (2024). This better represents the diversity of fields in the dataset, although in practice it does not generally make a big difference (see Figure 8).
  • When a model requires it, make sure to add the appropriate <bos> token at start of prompt (e.g., Gemma (Gemma Team et al., 2024)).
  • Restrict all inputs (with completions) to 2048 tokens for consistency across models. [^9] For current tasks this is only exhausted for a few MMLU instances.
  • Use the default model precision when evaluating (i.e., avoid options like load_in_8bit unless it produces identical results).
  • Other than the original instruction line for MMLU Hendrycks et al. (2021), we do not add any extra instructions. This is in view of previous work finding the subject information from instructions makes little changes to model ranking (Alzahrani et al., 2024), and to reduce additional sources of variation in the prompt.
Note that computational details, like batch size and type/state of GPU, can affect floating point operations such that answer choice decisions can flip if they are very close. This is hard to avoid unless one considers “ties” when answers are sufficiently close in confidence, we leave that for future consideration. A reference implementation of OLMES is released at https://github.com/allenai/olmes under the Apache 2.0 license.
@@ 4 OLMES: Summary and results
  • Use test set when available, otherwise validation. Sample 1000 instances if more than 1500 (Section 3.1)
@@
  • Follow recommendations for all other evaluation details (in Section 3.5)
Table 4 reports the overall, fully reproducible, OLMES scores for the 15 models across the benchmarks. An extended table with a total of 40 models is shown in Table 13 (Appendix D).
@@
OLMo-7B-0424 $66.9^{\dagger}$ $83.6^{\dagger}$ 85.9 $85.8^{\dagger}$ 80.1 $54.4^{\dagger}$ $68.6^{\dagger}$ 80.3 $76.1^{\dagger}$ 73.6 75.5
@@ 5 Related work
With model releases, performance on popular datasets is used to gauge the progress achieved e.g., OpenAI (2024) showing superhuman performance on benchmarks like MMLU. Such evaluation also guides community efforts towards understanding and sharing findings on what it takes to build a strong model (Touvron et al. (2023a, b); Biderman et al. (2023); Almazrouei et al. (2023); MosaicML (2023); Jiang et al. (2023); Gemma Team et al. (2024); Groeneveld et al. (2024) inter alia). However, given a model and a dataset, even for the frequently used datasets, there are varied practices in how accuracy on them is measured. Various work has shown that model evaluations are vulnerable to differences such as option position changes in multiple-choice questions (Zheng et al., 2024; Li et al., 2024), choice symbols, re-ordering of answer options, changing number of answer options (Wang et al., 2024), and task formulation (Alzahrani et al., 2024; Robinson et al., 2023; Khatun and Brown, 2024; Wiegreffe et al., 2023). Even minor formatting changes can cause large, generally arbitrary, score variations (Sclar et al., 2023).
The Holistic Evaluation of Language Models (HELM) benchmark (Liang et al., 2023), the Hugging Face Open LLM Leaderboard (Beeching et al., 2023), Mosaic Eval Gauntlet (Barton, 2024), Eleuther LM Evaluation Harness (Gao et al., 2023; Biderman et al., 2024), and Unitxt (Bandel et al., 2024) present efforts toward greater transparency and reproducibility of LLM evaluations. These frameworks generally describe and provide support for various task setups, presenting them as open choices to researchers and users. When specific default setups are given, the rationale is not always documented and thus not followed by others in subsequent work (see Tables 1 and  14).
@@ 6 Discussion
By identifying and reviewing common evaluation practices in the community, and performing experiments to resolve open questions, we present OLMES – an open, documented, reproducible, and practical evaluation standard. OLMES provides justified recommendations on decisions such as how to format dataset instances, the choice of in-context examples, task formulation, probability normalization, as well as other implementation details. The goal is for OLMES to be a useful guide for model developers to obtain signals as to whether their model is on track during training, and to compare final powerful base models. The practical choices encourage evaluations without unnecessary computation resources. The reproducible nature means that any evaluation done using OLMES can be directly compared to existing OLMES evaluations. We also document the rationales behind the choices made, guiding the community toward more justified evaluation practices. OLMES can be applied to current leaderboards and evaluation code bases to unify evaluation practices in the field.
@@
Future work and limitations. Future work includes adding more tasks to OLMES, covering tasks beyond MCQA such as generative tasks and chain-of-thought prompting. This will include standardizing how answers are extracted for evaluation, and for chat models how to split the prompt into messages. We welcome the community to contribute to OLMES, extending the principles of OLMES to new tasks.
@@
OLMES is a step towards standardizing LLM evaluations, ready to be incorporated into evaluation code bases for broad usage. OLMES facilitates robust and simplified comparisons of model performances, both for researchers during model training and development, and for developers in choosing models to build upon.
Acknowledgments